40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Thu Feb 24 10:44:06 2022, Radhika, DailyProgress, Cryo vacuum chamber, Double copper bar cooldown rerun started 02/23 IMG_3095.jpgIMG_3096.jpgIMG_3083.jpgIMG_3087.jpgIMG_3092.jpg
    Reply  Tue Mar 1 15:59:34 2022, Radhika, DailyProgress, Cryo vacuum chamber, Double copper bar cooldown rerun started 02/23 
       Reply  Fri Mar 4 13:18:09 2022, Radhika, DailyProgress, Cryo vacuum chamber, Double copper bar cooldown rerun started 02/23 02-24_cooldown_data.pdf02-24_cooldown_data_comparison.pdfcooldown_0224_bestfit.pdf
Message ID: 2724     Entry time: Thu Feb 24 10:44:06 2022     Reply to this: 2726
Author: Radhika 
Type: DailyProgress 
Category: Cryo vacuum chamber 
Subject: Double copper bar cooldown rerun started 02/23 

[Radhika, Aaron]

The goals for this cooldown are:

1) Securely re-attach the RTD to the cold head, to fix wacky readings.
2) Reduce thermal conductance between the outer shield and cold plate (/copper bar), to prevent the outer shield from cooling below 200K.
       2a) Stabilize the outer shield so that its contact area with the cold plate is consistent run-to-run. ([Note] Follow-up to an in-person comment from Rana: the positioning of the shields does not leave space for any clamps to secure the outer shield to the cold plate.)

On Tuesday 2/22, we opened up the bottom conflat of the T to check on the RTD spring-clamped to the cold head. I re-inserted the RTD and tightened the nut further than last run, and it seemed much more secure [Attachment 1]. I re-inserted the mylar "cap" covering the cold head [Attachment 2].

In the chamber body, we carefully passed the RTD leads through the inner shield and outer shield apertures to remove the outer shield. We did this without having to unclamp/remove the inner shield or any components inside, to preserve consistency with the last cooldown. A few pins were damaged in this process (from inner shield).

Once the outer shield was removed, we used kapton tape to secure strips of peek sheets to its bottom rim [Attachments 3,4]. The strips were taped at 2 points along the rim associated with the most wobble, with hopes of stabilizing the shield as much as possible.

On 2/23 I repaired the pins previously damaged. I also added kapton tape labels to the socket leads, corresponding to the shapes found on the RTD leads (semi-circle example in Attachment 5). This way it will be much easier to match the right pins and sockets in the future. 

I then bolted up the chamber (close-out pictures can be found on the QIL Google photo dump). The vacuum pump was turned on at 5:45pm, and the cryocooler was turned on at 7:08pm.

Attachment 1: IMG_3095.jpg  1.955 MB  | Hide | Hide all
IMG_3095.jpg
Attachment 2: IMG_3096.jpg  1.690 MB  | Hide | Hide all
IMG_3096.jpg
Attachment 3: IMG_3083.jpg  1.878 MB  | Hide | Hide all
IMG_3083.jpg
Attachment 4: IMG_3087.jpg  1.641 MB  | Hide | Hide all
IMG_3087.jpg
Attachment 5: IMG_3092.jpg  1.996 MB  | Hide | Hide all
IMG_3092.jpg
ELOG V3.1.3-